# Suppression of elliptic anisotropy inside jets: A new perspective for jet quenching

### v1

@@ II Elliptic anisotropies inside jets in elementary hadronic collisions
Another source of azimuthal-angle correlations is the dynamical process of the jet cascade. For example, in the parton shower stage, the parton splittings are more likely to happen at smaller angles Sjostrand:2014zea , which may enhances the probability to detect two particles near to each other in direction. In a typical configuration where there is only one large-angle splitting but much more small-angle splittings inside a jet, the jet may be roughly thought as two sub-jets. In this case, two particles from a same sub-jet will contribute to the correlations at a short-range (small differences in $\eta^{*}$), whereas those from different sub-jets may give rise to a back-to-back correlation, probably at a long-range. Besides the parton shower dynamics, the non-perturbative dynamics for the hadronization or fragmentation process may also affect the correlations to some extent. In addition, the interactions among the jet particles, at both partonic and hadronic stages, may have impacts on the azimuthal anisotropies. In particular, for the high-multiplicity jet production, in which the anisotropies arise from other sources become diminished, the rescatterings among the jet particles may lead to a hydrodynamics-like collective evolution, which may generate a long-range elliptic flow correlation CMS:2023iam ; Zhao:2024wqs .
@@
Figure 1 shows a comparison between the model results and the CMS data CMS:2023iam for the elliptic anisotropy coefficient $v_{2}$ inside jets in $pp$ collisions as a function of the charged-particle multiplicity inside jets. Here the anti-$k_{T}$ jets with $R=0.8$ Cacciari:2011ma are selected in the kinematic window $p_{T}>550$ GeV and $|\eta|<1.6$, and the $v_{2}\{2\}$ from two-particle correlations with a pseudorapidity gap $|\Delta\eta^{*}|>2$ are obtained using the charged particles inside jets with the cuts $|\eta|<2.4$, $p_{T}>0.3$ GeV and $3.00.3<j_{T}<$ 3.0 GeV CMS:2023iam . A Q-cumulant method Bilandzic:2010jr ; Zhao:2024wqs is used to calculate the $v_{2}$ in the model predictions.
Figure 2: (Color online) Elliptic anisotropies coefficient $v_{2}\{2,|\Delta\eta^{*}|>2.0\}$ as a function of jet transverse momentum $p_{T}^{\textrm{jet}}$ for inclusive jet production in $pp$ collisions at $\sqrt{s}=5.02$ TeV. Results for jets reconstructed at both hadronic (left panel) and partonic (right panel) levels calculated with three cuts on transverse momenta of jet particles, i.e. $p_{T}>0.3$, $p_{T}>1.0$ and $p_{T}>5.0$ GeV are shown respectively. Anti-$k_{T}$ jets with $R=0.8$ selected in region $|\eta|<1.6$ are used in calculations.
@@
Assuming that the PYTHIA8 model has captured the main features of the jet cascade dynamics relevant to the azimuthal correlations, we perform a calculation for the inclusive jet production in $pp$ collisions at $\sqrt{s}=5.02$ TeV and plot the results of the $v_{2}$ inside jets as a function of the transverse momentum of jet in Fig. 2. The left panel shows the results of $v_{2}\{2,|\Delta\eta^{*}|>2.0\}$ for the charged particles inside jets with three cuts on the particle transverse momentum, i.e. $p_{T}^{ch}>0.3$, $1.0$ and $5.0$ GeV, respectively. It is observed that in the studied regions the $v_{2}$ values are nontrivial, which slightly increase with the $p_{T}^{jet}$ for a certain $p_{T}^{ch}$ cut. At the same time, the $v_{2}$ obviously increase with an increasing $p_{T}^{ch}$ cut, which is partly due to that the harder particles which carry larger fractions of the jet momenta are more strongly limited by the momentum balance condition $\sum_{i}{\vec{j}_{T,i}}=0$.
@@ III Medium-induced modifications for inclusive jets in AA collisions
  T (GeV)   ||||   Prop. time (fm/c)   ||||    $\alpha_{s}$    ||||      Scatt.    $\langle v_{2}\rangle$
@@
Figure 3: (Color online) Results of $v_{2}\{2,|\Delta\eta^{*}|>2.0\}$ as a function of $p_{T}^{\textrm{jet}}$ for inclusive jet production in $pp$ and PbPb collisions at $\sqrt{s_{NN}}=5.02$ TeV, calculated with with two cuts on transverse momenta of jet particles as $p_{T}>0.3$ (top panel) and $p_{T}>5.0$ GeV (bottom panel), respectively. Anti-$k_{T}$ jets at partonic levels with $R=0.8$ selected in region $|\eta|<1.6$ are used in calculations. Results for two centrality classes of PbPb collisions, i.e., $0\!-\!10\%$ and $50\!-\!90\%$ are shown respectively.
@@
It is noted that, in the calculations with $p_{T}^{\textrm{parton}}>0.3$ GeV, the jet particles from the medium excitations, probably carrying small fractions of the jet momenta Luo:2018pto ; Wang:2016fds ; Chen:2017zte may contribute to the modification of $v_{2}$. For instance, with the parameter setting ’A’ in Tab. 1, $16\%$ of the jet particles are from the medium excitations in LBT model. As a test, we further calculate the $v_{2}$ with $p_{T}^{\textrm{parton}}>5$ GeV, where most of the excitation particles are excluded, and find similar suppression on $v_{2}$. For example, the $v_{2}$ calculated for $pp$ and for the medium with setting ’A’ are 0.316 and 0.259, respectively.
@@
Figure 4: (Color online) Trigger-particle-normalized two-particle azimuthal-angle distribution inside jets as a function of $\Delta\phi^{*}$ for $|\Delta\eta^{*}|>2.0$ in both $pp$ and PbPb ($0\!-\!10\%$) collisions.
@@
Figure 5: Jet-particle distributions in $\vec{j}_{T}$ plane for inclusive jet production in both $pp$ (left panel) and central PbPb (right panel) collisions at $\sqrt{s_{NN}}=5.02$ TeV. Anti-$k_{T}$ jets reconstructed at a partonic level with R=0.8 and $p_{T}^{jet}>100$ GeV are used. Distributions are obtained by averaging over $5\times 10^{5}$ collected jets with jet plane orientation in alignment.
@@
In the calculations for Fig. 5, the jet particles with $p_{T}^{\textrm{parton}}>0.3$ GeV are used to determine the jet plane, and for each generated jet system, the jet plane is rotated to be oriented along the $j_{x}$ direction Qiu:2011hf . By averaging over all the collected jets, one can observe in Fig. 5 that, the jets generated in $pp$ collisions exhibit an obvious elliptic distribution, whereas the quenched jets in AA collisions becomes more isotropic. The elliptic anisotropy for such distributions can be evaluated with $v_{2}=\langle\frac{j_{x}^{2}-j_{y}^{2}}{j_{x}^{2}+j_{y}^{2}}\rangle$ Ollitrault:1992bk ; Bilandzic:2010jr , which are found to be 0.33 and 0.16 for $pp$ and AA, respectively. Moreover, a double-peak structure can be found in both of the two distributions, indicating that such a back-to-back configuration plays an important role in the generation of elliptic anisotropy inside jets.

### v2

@@ II Elliptic anisotropies inside jets in elementary hadronic collisions
The various dynamical processes inherent in jet evolution and formation constitute another source of azimuthal-angle correlations. For example, in the parton shower stage, the parton splittings are more likely to happen at smaller angles Sjostrand:2014zea , which may enhances the probability to detect two particles near to each other in direction. In a typical configuration where there is only one large-angle splitting but much more small-angle splittings inside a jet, the jet may be roughly thought as two sub-jets. In this case, two particles from a same sub-jet will contribute to the correlations at a short-range (small differences in $\eta^{*}$), whereas those from different sub-jets may give rise to a back-to-back correlation, probably at a long-range. Besides the parton shower dynamics, the non-perturbative dynamics for the hadronization or fragmentation process may also affect the correlations to some extent. In addition, the interactions among the jet particles, at both partonic and hadronic stages, may have impacts on the azimuthal anisotropies. In particular, for the high-multiplicity jet production, in which the anisotropies arise from other sources become diminished, the rescatterings among the jet particles may lead to a hydrodynamics-like collective evolution, which may generate a long-range elliptic flow correlation CMS:2023iam ; Zhao:2024wqs .
@@
[…]
@@
Given PYTHIA8’s demonstrated ability to reasonably describe the $v_{2}$ data across most multiplicity regimes (excluding high multiplicity regions), we perform a calculation for the inclusive jet production in $pp$ collisions at $\sqrt{s}=5.02$ TeV and plot the results of the $v_{2}$ inside jets as a function of the transverse momentum of jet in Fig. 2. The left panel shows the results of $v_{2}\{2,|\Delta\eta^{*}|>2.0\}$ for the charged particles inside jets with three cuts on the particle transverse momentum, i.e. $p_{T}^{ch}>0.3$, 1.01.01.0 and 5.05.05.0 GeV, respectively. It is observed that in the studied regions the $v_{2}$ values are nontrivial, which slightly increase with the $p_{T}^{jet}$ for a certain $p_{T}^{ch}$ cut. At the same time, the $v_{2}$ obviously increase with an increasing $p_{T}^{ch}$ cut, which is partly due to that the harder particles which carry larger fractions of the jet momenta are more strongly limited by the momentum balance condition $\sum_{i}{\vec{j}_{T,i}}=0$.
@@ III Medium-induced modifications for inclusive jets in AA collisions
  T (GeV)   |||   Prop. time (fm/c)   |||    $\alpha_{s}$    |||      Scatt.    $\langle v_{2}\rangle$
@@
[…]
@@
[…]
@@
Figure 4: (Color online) Trigger-particle-normalized two-particle azimuthal-angle distribution inside jets as a function of $\Delta\phi^{*}$ for $|\Delta\eta^{*}|>2.0$ in both $pp$ and PbPb ($0\%$ 10 %) collisions.
@@
[…]
@@
In the calculations for Fig. 5, the jet particles with $p_{T}^{\textrm{parton}}>0.3$ GeV are used to determine the jet plane, and for each generated jet system, the jet plane is rotated to be oriented along the $j_{x}$ direction Qiu:2011hf . By averaging over all the collected jets, one can observe in Fig. 5 that, the jets generated in $pp$ collisions exhibit an obvious elliptic distribution, whereas the quenched jets in AA collisions becomes more isotropic. The elliptic anisotropy for such distributions can be evaluated with $v_{2}=\langle\frac{j_{x}^{2}-j_{y}^{2}}{j_{x}^{2}+j_{y}^{2}}\rangle$ Ollitrault:1992bk ; Bilandzic:2010jr , which are found to be 0.33 and 0.16 for $pp$ and AA, respectively. Moreover, a double-peak structure can be found in both of the two distributions, indicating that such a back-to-back configuration plays an important role in the generation of elliptic anisotropy inside jets. We observe that the values of $v_{2}$ calculated with such a method are larger than the results of $v_{2}\{2,|\Delta\eta^{*}|>2.0\}$ obtained with the two-particle correlation method (Fig. 3). This discrepancy can be partly attributed to the small number of particles and the less preserved longitudinal symmetry in the jet systems. Notably, despite these methodological differences, both approaches consistently predict the medium-induced suppression of $v_{2}$.